OpenStudy (anonymous):

What is the midpoint of (–2, 0) and (4, –2)? (0, 0) (2, –2) (1, –1) (–6, 2)

OpenStudy (jamesj):

Well, the x-ordinate has to be midway between -2 and 4. So that already gives you a big clue as to which is the right answer

OpenStudy (anonymous):

my guess was b. just wanted to make sure.

OpenStudy (shubhamsrg):

Really ? What makes you say B ?

OpenStudy (jamesj):

What is the x that is midway between -2 and 4? It is \[ \frac{-2 + 4}{2} = \frac{2}{2} = ... \]